A stone is thrown vertically upwards. When particle is at a height half of its maximum height, its speed is 10m/s; then the maximum height attained by the particle is: [g = 10m/s2]


        (a)    8 m


        (b)    10 m


        (c)    15 m


        (d)    20 m


Grade:10

2 Answers

Jitender Pal
10 years ago

(b)
Let h be maximum height to be attained by stone.
Velocity with which particle is thrown
    u= √2gh
    v^2=u^2+2gh
    when H=h/2,u= √2gh,v=10 m/s   g=-g
    v^2-u^2=  2gH 

2gh-100=  2g h/2
        2 g  h/2  =100 

   h =10 m.

Arpit Aggarwal
4 years ago
V^2 - u^2 = 2aS
10^2 - u^2 = -2gH/2
-gH = 100 - u^2
-gu^2 / (2g) = 100 - u^2
u^2 / 2 = u^2 - 100
100 = u^2 / 2
u^2 = 200
H = u^2 / (2g)
= 200 / (2 × 10)
= 10 m
Maximum height attained by it is 10 m